OpenStudy (anonymous):

A convenience store currently stocks 48 bottles of mineral water. The bottles have two sizes of either 20 or 40 ounces each. The average volume per bottle the store currently has in stock is 35 ounces. How many 40 ounce bottles are in stock?

OpenStudy (aaronandyson):

Well.. "(20x+40y)/(x+y)=35 x+y=48 x=48-y Use this value of x in the first equation (960-20y+40y)/(48-y+y)=35 (960+20y)/48=35 (multiply both sides by 48) 960+20y=1680 (subtract 960 from both sides) 20y=720 (divide both sides by 20) y=36 so there are 36 40oz bottles "

OpenStudy (anonymous):

where did the 960 come from?

OpenStudy (aaronandyson):

20*48 = 960

OpenStudy (anonymous):

got it thanks

OpenStudy (aaronandyson):

YW ;)